OpenStudy (anonymous):

Find the exact value of each expression. sin 36° cos 9° + cos 36° sin 9°

OpenStudy (anonymous):

sin(36+9)=sin(45)

OpenStudy (anonymous):

\[ \sin(45)=\frac{\sqrt 2}{2} \]

OpenStudy (anonymous):

sin(a+b)= sin(a)cos(b) + sin(b)cos(a)

OpenStudy (anonymous):

Thank you! :)